# NaP2: A Benchmark for Naturalness and Privacy-Preserving Text Rewriting by Learning from Human

### v1

@@ Abstract
Increasing concerns about privacy leakage issues in academia and industry arise when employing NLP models from third-party providers to process sensitive texts. To protect privacy before sending sensitive data to those models, we suggest sanitizing sensitive text using two common strategies used by humans: i) deleting sensitive expressions, and ii) obscuring sensitive details by abstracting them. To explore the issues and develop a tool for text rewriting, we curate the first corpus, coined NaP2, through both crowdsourcing and the use of large language models (LLMs). Compared to the prior works based on differential privacy, which lead to a sharp drop in information utility and unnatural texts, the human-inspired approaches result in more natural rewrites and offer an improved balance between privacy protection and data utility, as demonstrated by our extensive experiments.
@@ 1 Introduction
Data sharing and information dissemination between AI models are pivotal in the AI era, particularly since the emergence of Large language models (LLMs). The remarkable performance of LLMs benefit from a large amount of shared and publicly available data. However, it is still challenging to balance between data privacy and information utility when training and utilizing such LLMs Pan et al. (2020). Users or downstream applications often interact with commercial LLMs by directly inputting raw text. Such interactions can inadvertently expose sensitive data, such as personally identifiable information (PII), to untrusted service or LLM providers Utpala et al. (2023).
Redaction and anonymization techniques are widely applied to remove PII from texts, but they suffer from three major drawbacks Sánchez et al. (2014). First, after anonymization, mentions of PII are either redacted or replaced by their entity types so that processed texts become unnatural. Downstream applications need to be adapted or fine-tuned to cope with such unnatural texts. Second, it is still possible to recover private attributes from PII scrubbed text via reasoning Mireshghallah et al. (2023); Staab et al. (2023). Third, the presence of blacked-out parts or entity types may raise the awareness of a document’s sensitivity in front of potential attackers.
ORI: I am Cindy .
I am recovering from
ulnar nerve surgery .
PER: I just had surgery.
@@
DEL: I am Cindy.
OBS: Hi, I am Cindy, I am just
recovered from medical treatment
T5 trained on NaP2:
Output: I am recovering from llness.
DP method
DPNR: YYYYYYYYY.
DP-Forward: My name is Cindy and I am
recovering from ulnar nerve surgery.
@@
Table 1: An example of rewriting a text (ORI) using deleting (DEL) and obscuring (OBS) as the strategies based on personal information (PER).
@@
Alternatively, differential privacy (DP) provides a theoretical privacy guarantee for data release or dissemination mechanisms Dwork (2006). Prior works sanitize texts by perturbing texts either at the word-level or the sentence-level Mattern et al. (2022); Igamberdiev and Habernal (2023); Igamberdiev et al. (2022a). In order to reach a high-level of privacy guarantee, substantial noise need to injected into texts or their representations so that information utility drops sharply and the meanings of texts are changed significantly (see Table 1). Therefore, determining an optimal trade-off between privacy and utility for data release remains an unresolved challenge.
@@
To address the shortcoming of prior methods, we propose an approach that adopts human text editing strategy inspired by  Strengers et al. (2020), specifically deleting and obscuring to improve the naturalness and utility of rewritten texts while ensuring privacy. As shown in Table 1, given an utterance involving personal information stated in a persona, the strategy deleting simply removes all words mentioning sensitive information from the utterance, while obscuring substitutes sensitive expressions for more abstract and general expressions. Both strategies aim to make rewritten texts as natural as possible such that i) they do not raise the awareness of potential attackers that rewrites are sanitized; and ii) downstream applications can directly process such natural rewrites without fine-tuning their models for any unnatural parts of texts.
@@
To evaluate strategy-specific rewriting models, we construct the first Naturalness and Privacy Preserving Rewriting corpus, coined NaP2, based on the open-domain dialogue corpus PERSONA-CHAT Zhang et al. (2018). We recruit university students to manually rewrite 895 utterances involving personal information as the manual evaluation set. To promote the development of diverse open-source solutions for this task, we apply GPT4 to generate 3900 synthetic examples as the synthetic training set because GPT4 demonstrates the best performance on PERSONA-CHAT among all evaluated models. We also design multiple automatic and human evaluation metrics for this task, including a novel privacy metric Privacy_NLI. It utilizes a Natural Language Inference (NLI) model Liu et al. (2019) to determine if a rewrite entails a personal information or not. The extensive comparative studies between the models trained on our corpus and the state-of-the-art (SOTA) text sanitization methods demonstrate the underlying challenges and yield the following key findings:
@@
  • The privacy metric Privacy_NLI aligns well with the human judgements by having a Spearman’s ranking correlation of 0.70.
@@
2 Preliminary
As our task is closely related to local differential privacy (LDP) Xiong et al. (2020), this section mainly introduces the preliminary concepts of LDP and Context-Aware LDP Acharya et al. (2020).
LDP.
For private statistical data publication, LDP provides strong and provable privacy preservation without assuming that data collectors can be trusted. In this setting, each participant locally perturbs her/his private data with a randomized mechanism and transfers the perturbed output to a data collector. The data collector acquires statistical information from the perturbed data received from all participants without compromising the individual’s privacy.
Definition 2.1 (Pure Local Differential Privacy Kasiviswanathan et al. (2011)).
Let $\epsilon\geq 0$, a randomized algorithm $Q:\mathcal{X}\rightarrow\mathcal{Y}$ is $\epsilon-$locally differential privacy, if for all $x,x^{\prime}\in\mathcal{X}$ and $y\in\mathcal{Y}$,
$\frac{Q(y|x)}{Q(y|x^{\prime})}<e^{\epsilon}$
where $Q(y|x)$ can also be viewed as a conditional distribution. If the privacy budget $\epsilon$ is small or zero, it is difficult or infeasible to distinguish between $x$ and $x^{\prime}$ based on the outputs of $Q$. However, it is questionable if such perturbed data is still useful for data analysts or downstream applications (Mattern et al., 2022).
This definition assumes that all elements in $x$ are equally sensitive and all $x$ share the same privacy constraint $e^{\epsilon}$ regardless of how different $x$ and $x^{\prime}$ are. For NLP applications, Mattern et al. (2022) observe that a tight universal privacy budget leads to substantial grammatical errors produced by word-level DP mechanisms, while a high budget easily compromises individual privacy. Therefore, the privacy budget of $x$ and $x^{\prime}$ should vary depending on the semantic differences between $x$ and $x^{\prime}$ or whether $x$ or $x^{\prime}$ contains sensitive information.
To address the limitations, Acharya et al. (2020) propose context-aware LDP that employs a different privacy constraint based on differences between $x$ and $x^{\prime}$.
Definition 2.2 (Context-Aware LDP Acharya et al. (2020)).
Let ${\mathbf{E}}\in\mathbb{R}^{m\times m}$ be a matrix of non-negative entries and $\epsilon_{x,x^{\prime}}$ denote the $(x,x^{\prime})$th entry of ${\mathbf{E}}$, a random algorithm $Q$ is ${\mathbf{E}}-$LDP, if for all $x,x^{\prime}\in\mathcal{X}$ and $y\in\mathcal{Y}$,
$\frac{Q(y|x)}{Q(y|x^{\prime})}<e^{\epsilon_{x,x^{\prime}}}$
The matrix ${\mathbf{E}}$ can be constructed by using different functions. Metric-based LDP Alvim et al. (2018) can also be viewed as a special case of context-aware LDP by requiring $\epsilon_{x,x^{\prime}}=\epsilon d(x,x^{\prime})$, where $d(x,x^{\prime})$ is the metric between $x$ and $x^{\prime}$.
@@ 3.1 Problem Definition
Given an utterance ${\bm{x}}$ and a sentence ${\bm{p}}$ describing personal information, the task of naturalness and privacy-preserving rewriting aims to map ${\bm{x}}$ into a natural sentence ${\bm{y}}$ such that ${\bm{y}}\in\mathcal{Y}^{n}$ does not reveal the personal information in ${\bm{p}}$ and maximally preserves the non-private content in ${\bm{x}}$. We define a natural sentence as one that is grammatically correct, fluent, and does not contain any artifacts such as blacked-out words or special symbols indicating omitted sensitive information. The rewrite space $\mathcal{Y}^{n}$ contains only natural sentences. Compared with differential privacy (DP) mechanisms that prevent privacy leakage during model training Abadi et al. (2016a), this task focuses on privacy-preserving data publishing or privacy protection at inference time.
@@
Relation to LDP. A probabilistic rewriting model can be viewed as a random mechanism $Q(y|x)$ that maps an input text $x$ into a word sequence $y$ inside a constraint output space $\mathcal{Y}$, which only contains natural texts. Given a pair of semantically similar texts $x$ and $x^{\prime}$, where only $x$ contains sensitive information, a rewriting model $Q(y|x)$ implements metric-based LDP by enforcing the distribution divergence in log-scale between $Q(y|x)$ and $Q(y^{\prime}|x^{\prime})$ to be smaller than $\epsilon d(x,x^{\prime})$. As a result, perturbed texts are more similar than their original counterparts because mentions of private information are either removed or obscured.
Corpus Overview. Our corpus NaP2 consists of a small manually curated dataset for both training and testing (Sec. 3.2), and a large synthetic dataset distilled from GPT-3.5 turbo and GPT4 for training data augmentation (Sec. 3.3). According to our evaluation stated below, human writes with obscuring achieve the best trade-off between privacy and utility, and the naturalness of GPT4 generated texts is on par with that of human rewrites.
@@ 3.2 Manually Curated Corpus
The corpus PERSONA-CHAT associates each multi-turn chit-chat with two personas, each of which is a set of sentences describing the corresponding personality. Hence, it is straightforward to measure if an utterance leaks personal information in the relevant persona. From another point of view, a persona can be regarded as a user-specific privacy profile, which states what information needs to be protected. For instance, one user might consider their marital status as sensitive information requiring privacy protection, while another user may not prioritize it.
The manual created evaluation set extends the test set of PERSONA-CHAT with human-authored rewrites. As not all utterances reveal private information in personas, we apply the automatic alignment methods to pair an utterance involving personal information with the corresponding sentence in a persona.
Formally, given a dialogue $\mathcal{D}$, suppose there are $m$ utterances $\mathcal{X}_{i}=\{{\bm{x}}_{1},{\bm{x}}_{2},...,{\bm{x}}_{m}\}$ associated with a persona $\mathcal{P}_{i}=\{{\bm{p}}_{1},{\bm{p}}_{2},...,{\bm{p}}_{n}\}\}$, we aim to compute an alignment score $s_{ij}$ between ${\bm{x}}_{i}\in\mathcal{X}_{i}$ and ${\bm{p}}_{j}\in\mathcal{P}_{i}$ indicating to what degree ${\bm{x}}_{i}$ leaks personal information in ${\bm{p}}_{j}$.
@@
In our preliminary experiments, we observe that even though annotators endeavor to generate decent rewrites, many of them could not clearly identify and strictly stick to the required strategies. Therefore, we prepare a small sample of pairs as a pre-test to select qualified annotators. In addition, we employ a rigorous procedure for quality check. In particular, we wrap up 15 sentence-persona pairs as a batch and ask annotators to rewrite them using the required strategies. Then, we manually check the rewritten batches, we only accept those that are written using the required strategy. The averaged acceptance rate of the rewrites is 47.97%, demonstrating the challenge of collecting a high-quality rewriting dataset with specific rewriting requirements. As a result, we collect 895 pairs annotated with one rewrite per strategy. We further split this corpus into a cross-validation (CV) set, a validation and a hold-out test set with 655, 140 and 100 instances, respectively.
@@ 3.3 Synthetic Data Augmentation
Rewrite this sentence, $<$deleting / obscuring$>$ any private information. Example rewrites are: <\$absentcurrency-dollar<\\$< \$$IN-CONTEXT\_EXAMPLES$$>$ Only return the rewritten sentence, nothing else. Private information present is: [\$PERSONA]. Sentence to rewrite is: [\$UTTERANCE].
@@ 3.4 Human Evaluation
Three university students are recruited to check their quality on a set of 100 instances sampled from the test set of the manual corpus. Hence, an utterance-persona pair in the sample includes a human rewrite, a rewrite from GPT-3.5 turbo and GPT4 respectively. For each rewrite, a student is instructed to answer the following questions from the perspectives of privacy leakage (Q1), semantic relevance (Q2) and naturalness (Q3) which is detailed in Appendix A.1.
Each question is answered by three university students. To deal with possible disagreements, we take the majority vote as the final answer.
@@ 4.1 Rewriting Models
We compare the SOTA privacy-preserving rewriting models DPNR Lyu et al. (2020), DP-Forward Du et al. (2023), and the zero-shot LLMs with the T5-Base models fine-tuned on our corpus, with or without synthetic data augmentation. The word-level DP method DPNR and the sentence-level DP model DP-Forward are fine-tuned based on a T5-Base model, which is pre-trained to map inputs to their outputs. All implementation details can be found in Appendix A.3. DPNR. It stands for Differentially Private Neural Representation, which applies Laplace noise to distributed representations of words in order to randomly drop sensitive words or replace sensitive words with non-sensitive ones. We compare the cosine similarity between each word in an input utterance with those in the corresponding persona, and pick the top-$k$ most similar ones.
@@
Llama-13B_zeroshot-obscuring 79.28% 16.00% 40.86% 40.12%
Llama-13B_zeroshot-deleting 77.14% 14.00% 68.28% 67.53%
@@
Llama-paraph. Mattern et al. (2022) points out the limitations of word-level LDP and propose to paraphrase input texts with lower temperature to achieve a sentence-level LDP. We implement this approach by using Llama-13B.
DP-Prompt.Utpala et al. (2023) utilizes zero-shot prompting and large language model to generate document paraphrasing to prevent author de-anonymization attack which comprise the privacy of text owner.
DP-BART. The method is a privatized text rewriting system incorporates LDP. The system leverages the LPD paradgram to perform model rewriting using BART model to protect input data which tackles same challenge like us.
FLAIR-SCRUBBING. we also adapt the scrubbing method used in  Lukas et al. (2023) as our baseline. We employ FLAIR-SCRUBBING as our method to test if this automatics method can effectively remove private information from sentence.
Zero-Shot LLMs. To compare with the LLMs fine-tuned on our corpus, we apply the same prompts to the same pre-trained LLMs without any training. Specifically, we consider T5-Base, Llama-13B, GPT-3.5 turbo and GPT4 and apply the prompt template introduced in Sec. 3.3. To distinguish from the fine-tuned models, the T5-Base and Llama-13B in the zero-shot setting is referred to as T5_zeroshot and Llama-13B_zeroshot, respectively.
T5-NaP2. By using the same prompts as the zero-shot version, we fine tune T5-Base on the training set of the manually curated corpus, with or without augmenting them with synthetic data. The prompts are similar to those used by zero-shot models detailed in A.2.

### v2

@@ Abstract
The widespread use of cloud-based Large Language Models (LLMs) has heightened concerns over user privacy, as sensitive information may be inadvertently exposed during interactions with these services. To protect privacy before sending sensitive data to those models, we suggest sanitizing sensitive text using two common strategies used by humans: i) deleting sensitive expressions, and ii) obscuring sensitive details by abstracting them. To explore the issues and develop a tool for text rewriting, we curate the first corpus, coined NaP2, through both crowdsourcing and the use of large language models (LLMs). Compared to the prior works on anonymization, the human-inspired approaches result in more natural rewrites and offer an improved balance between privacy protection and data utility, as demonstrated by our extensive experiments. Researchers interested in accessing the dataset are encouraged to contact the first or corresponding author via email.
@@ 1 Introduction
Data sharing and information dissemination between AI models are pivotal in the AI era, particularly since the emergence of Large language models (LLMs). The remarkable performance of LLMs benefits from a large amount of shared and publicly available data. However, it is still challenging to balance between data privacy and information utility when training and utilizing such LLMs Pan et al. (2020) with massive amount of data. Users or applications often interact with commercial LLMs by directly inputting raw text. Such interactions can inadvertently expose sensitive data, such as personally identifiable information (PII), to untrusted service or LLM providers Utpala et al. (2023).
Redaction and anonymization techniques are widely applied to remove PII from texts, but they suffer from three major drawbacks Sánchez et al. (2014). First, after anonymization, mentions of PII are either redacted or replaced by their entity types so that processed texts become unnatural as it breaks grammatical flow, coherence and semantic clarity of sentences. Downstream applications need to be adapted or fine-tuned to cope with such unnatural texts. Second, it is still possible to recover private attributes from PII scrubbed text via reasoning Mireshghallah et al. (2023); Staab et al. (2023). Third, the presence of blacked-out parts or entity types may raise the awareness of a document’s sensitivity in front of potential attackers.
ORI: I have two teenage boys.
I have been to Los Angeles
a few years ago.
PER: I am a single mom of two boys.
@@
DEL: I have been to Los Angeles
a few years ago.
OBS: I have some children.
I have been to Los Angeles
a few years ago.
T5-Base trained on NaP2:
Output: I have been to Los Angeles
a few years ago.
FLAIR-SCRUBBING:
Output: I have <MASK> teenage boys.
I have been to <MASK> <MASK>.
DP-Prompt:
$\epsilon$-10: Junior
$\epsilon$-100: I have two teenage boys.
I have been to Los Angeles
a few years ago.
@@
Table 1: An example of rewriting a text (ORI) using deleting (DEL) and obscuring (OBS) as strategies based on personal information (PER). Output shows the T5-Base model finetuned with NaP2. Also shown are results from FLAIR-SCRUBBING and DP-Prompt using $\epsilon$-10 and $\epsilon$-100.
The recent work on text anonymization Dou et al. (2024) introduced the task of self-disclosure abstraction, which involves rephrasing sensitive information into less specific terms while preserving utility (e.g., "I’m 16F" to "I’m a teenage girl"). A user study showed that 82% of participants responded positively to the system, underscoring its practical relevance. However, the study focuses exclusively on rewriting mentions of private attributes, and the accompanying corpus includes only annotated spans of private attributes without human-authored reference rewrites, limiting its suitability for reference-based evaluation metrics.
@@
Alternatively, differential privacy (DP) provides a theoretical privacy guarantee for data release or dissemination mechanisms Dwork (2006). Prior works sanitize texts by perturbing texts either at the word-level or the sentence-level Mattern et al. (2022); Igamberdiev and Habernal (2023); Igamberdiev et al. (2022a). In order to reach a high-level of privacy guarantee, substantial noise needs to be injected into texts or their representations so that information utility drops sharply and the meanings of texts are changed significantly (see Table. 1). Therefore, determining an optimal trade-off between privacy and utility for data release remains an unresolved challenge.
@@
To address limitations of prior methods, we propose a human-inspired text editing approach—drawing on deleting and obscuring strategies Strengers et al. (2020)—to enhance the naturalness and utility of rewritten texts while ensuring privacy, aligning with the suppression and generalization principles of k-anonymity Sweeney (2002) originally developed for structured data. As shown in Table 1, given an utterance involving personal information stated in a persona, the strategy deleting simply removes all words mentioning sensitive information from the utterance, while obscuring substitutes sensitive expressions for more abstract and general expressions. In our example, the user requires explicitly rewrite private information about "a single mom of two boys". deleting removes entire part about this information and remain other parts untouched. While obscuring obscures the information about "boys" and "teenager" to simply "children" which generalizes the information to be protected. Both strategies aim to make rewritten texts as natural as possible such that i) they do not raise the awareness of potential attackers that rewrites are sanitized; and ii) downstream applications can directly process such natural rewrites without fine-tuning their models for any unnatural parts of texts.
To evaluate strategy-specific rewriting models, we construct the first Naturalness and Privacy Preserving Rewriting corpus, coined NaP2, based on the open-domain dialogue corpus PERSONA-CHAT Zhang et al. (2018). Unlike prior work that focuses solely on private attributes, our corpus incorporates text-based personalized privacy profiles. Hence, detection of personal information cannot be formulated as a multi-class classification task. We recruit university students to manually rewrite 895 utterances involving personal information as the manual evaluation set.
@@
To promote the development of diverse open-source solutions for this task, we apply GPT4 to generate 3900 synthetic examples as the synthetic training set because GPT4 demonstrates the best performance on PERSONA-CHAT among all evaluated models. We also design multiple automatic and human evaluation metrics for this task, including a novel privacy metric Privacy_NLI. It utilizes a Natural Language Inference (NLI) model Liu et al. (2019) to determine if a rewrite entails a personal information or not. The extensive comparative studies between the models trained on our corpus and the state-of-the-art (SOTA) text sanitization methods demonstrate the underlying challenges and yield the following key findings:
@@
  • The privacy metric Privacy_NLI aligns well with the human judgments by having a Spearman’s ranking correlation of 0.70.
@@ 2.1 Problem Definition
Given an utterance ${\bm{x}}$ and a sentence ${\bm{p}}$ describing personal information, the task of naturalness and privacy-preserving rewriting aims to map ${\bm{x}}$ into a natural sentence ${\bm{y}}$ such that ${\bm{y}}\in\mathcal{Y}^{n}$ does not reveal the personal information in ${\bm{p}}$ and maximally preserves the non-private content in ${\bm{x}}$. We define a natural sentence as one that is grammatically correct, fluent, and does not contain any artifacts such as blacked-out words or special symbols indicating omitted sensitive information. The rewrite space $\mathcal{Y}^{n}$ contains only natural sentences with maximum sequence length of $n$. Compared with DP mechanisms that prevent privacy leakage during model training Abadi et al. (2016a), this task focuses on privacy-preserving data publishing or privacy protection at inference time while uploading user query.
@@
Dataset Year Source Human Synthetic Rewrite Type Form of Private info Size
NAP2 (Ours) 2025 PERSONACHAT Delete / Obscure privacy profile Small
Self-Disclosure Dou et al. (2024) 2024 Reddit Post Obscure text span Small
SythPAI Yukhymenko et al. (2024) 2024 Reddit sytle PII Large
Text Anon. Benchmark (TAB) Pilán et al. (2022) 2022 ECHR legal cases Masking PII Large
TextWashKleinberg et al. (2022) 2022 Wikipedia Bio Masking PII Medium
Table 2: Comparison of recent datasets for text anonymization or privacy-preserving rewriting. Human and Synthetic refers to if there is any generated rewrite for the dataset either from human or LLMs. Rewrite type includes deletion, obfuscation. Masking indicates the detection private entities is redacted as entity type
Corpus Overview.
Our corpus NaP2 consists of a small manually curated dataset for both training and testing (Sec. 2.2), and a large synthetic dataset distilled from GPT-3.5 turbo and GPT4 for training data augmentation (Sec. 2.3). According to our evaluation stated below, human rewrites with obscuring achieve the best trade-off between privacy and utility, and the naturalness of GPT4 generated texts is on par with that of human rewrites.
Comparison with existing datasets
Our dataset stands out from other recent anonymization datasets by providing both human and synthetic rewrites based on diversified privacy profiles, rather than simple PII spans. Unlike datasets such as Self-Disclosure, which rely on LLM-generated rewrites focused on detected text spans, NaP2 introduces more explicit rewrite operations including deletion and obscuration. Additionally, while other datasets emphasize masking or detection, NaP2 offers more natural rewrites grounded in persona-level privacy, the rewrite option are clearly stated as obscure and delete which can facilitate different privacy protection level.
@@ 2.2 Manually Curated Corpus
The corpus PERSONA-CHAT associates each multi-turn chit-chat with two personas, each of which is a set of sentences describing the corresponding personality. Detailed information for PERSONA-CHAT are displayed in the Appendix. A.2 Hence, it is straightforward to measure if an utterance leaks personal information in the relevant persona. From another point of view, a persona can be regarded as a user-specific privacy profile, which states what information needs to be protected. For instance, one user might consider their marital status as sensitive information requiring privacy protection, while another user may not prioritize it.
The manual created evaluation set extends the test set of PERSONA-CHAT with human-authored rewrites. As not all utterances reveal private information in personas, we apply the automatic alignment methods to pair an utterance involving personal information with the corresponding sentence in a persona. Formally, given a dialogue $\mathcal{D}$, suppose there are $m$ utterances $\mathcal{X}_{i}=\{{\bm{x}}_{1},{\bm{x}}_{2},...,{\bm{x}}_{m}\}$ associated with a persona $\mathcal{P}_{i}=\{{\bm{p}}_{1},{\bm{p}}_{2},...,{\bm{p}}_{n}\}$, we aim to compute an alignment score $s_{ij}$ between ${\bm{x}}_{i}\in\mathcal{X}_{i}$ and ${\bm{p}}_{j}\in\mathcal{P}_{i}$ indicating to what degree ${\bm{x}}_{i}$ leaks personal information in ${\bm{p}}_{j}$.
@@
In our preliminary experiments, we observe that even though annotators endeavor to generate decent rewrites, many of them could not clearly identify and strictly stick to the required strategies. Therefore, we prepare a small sample of pairs as a pre-test to select qualified annotators. In addition, we employ a rigorous procedure for quality check. We wrap up 15 sentence-persona pairs as a batch and ask annotators to rewrite them using the required strategies. Then, we manually check the rewritten batches, we only accept those that are written using the required strategy. The averaged acceptance rate of the rewrites is 47.97%, demonstrating the challenge of collecting a high-quality rewriting dataset with specific rewriting requirements. As a result, we collect 895 pairs annotated with one rewrite per strategy. We further split this corpus into a cross-validation (CV) set, a validation and a hold-out test set with 655, 140 and 100 instances, respectively.
@@ 2.3 Synthetic Data Augmentation
Rewrite this sentence, <deleting / obscuring> any private information. Example rewrites are: <IN-CONTEXT_EXAMPLES> Only return the rewritten sentence, nothing else. Private information present is: [\$PERSONA]. The sentence to rewrite is: [\$UTTERANCE].
@@ 2.4 Human Evaluation
Three university students are recruited to check their quality on a set of 100 instances sampled from the test set of the manual corpus. Hence, an utterance-persona pair in the sample includes a human rewrite, a rewrite from GPT-3.5 turbo and GPT4 respectively. For each rewrite, a student is instructed to answer the following questions from the perspectives of privacy leakage (Q1), semantic relevance (Q2) and naturalness (Q3) which is detailed in Appendix A.1. Each question is answered by three university students. To deal with possible disagreements, we take the majority vote as the final answer. For annotation, the three annotators achieved Fleiss’ KappaFalotico and Quatto (2015) inter-annotator agreement score with 0.47 which is acceptable for classification problem.
@@ 3.1 Rewriting Models
In this section, we establish a baseline approach to assess the efficacy of existing privacy protection solutions in removing private content within textual messages. This evaluation is pivotal to addressing the critical question: "Are current privacy protection solutions adequately equipped to conceal privacy-sensitive content in utterances?" We compare the SOTA privacy rewriting models
To have a comprehensive comparison of privacy-preserving methods during model inference time, we compare several state-of-the-art privacy-preserving rewriting models, including DPNR Lyu et al. (2020), DP-Forward Du et al. (2023), Llama-paraph, DP-Prompt, FLAIR-SCRUBBING, and DP-BART. DPNR applies Laplace noise to word representations, while DP-Forward perturbs embedding matrices for sentence-level LDP detailed explaination for each baseline method can be foud in Appendix. A.4.
To assess the rewriting ability of models in zero shot, we consider LLMs fine-tuned on our corpus and apply the same prompts to the same pre-trained LLMs without any additional training. Specifically, we consider T5-Base, Llama2-13B, GPT-3.5 turbo, and GPT4, and apply the prompt template introduced in Sec. 2.3. To distinguish them from the fine-tuned models, T5-Base and Llama2-13B in the zero-shot setting are referred to as T5_zeroshot and Llama2-13B_zeroshot, respectively. In order to test our dataset and synthetic data augmentation, we consider T5-Base fine-tuned on our corpus, with or without synthetic data augmentation. We further use BART to fine-tune our dataset to emphasize the effectiveness of our datasets. In addition, we assess T5-NaP2 fine-tuned with and without DP-SGD Abadi et al. (2016a) to evaluate the impact of differential privacy on model performance.
@@
Llama2-13B_zeroshot-obscuring 79.28% 16.00% 40.86% 40.12%
Llama2-13B_zeroshot-deleting 77.14% 14.00% 68.28% 67.53%